# Principles and Guidelines for Evaluating Social Robot Navigation Algorithms

### v1

@@ Abstract
Figure 1: We identify eight broad principles of social robot navigation - including safety, comfort, legibility, politeness, social competency, agent understanding, proactivity, and contextual appropriateness - which serve to motivate specific guidelines for experiments, metrics, scenarios, benchmarks, datasets, and simulators. Principles and guidelines are labeled with two-letter codes, with P for principles, R for real-world issues, M for metrics, N for scenarios, B for benchmarks, D for datasets, and S for simulators.
@@ I Introduction
The study of social robot navigation has a long history, but a crisp definition of what makes navigation “social” remains elusive. Researchers on social robot navigation often have a personal sense of what it is and use that intuition to guide their research into how to make robots move better around people, but the field does not yet have a consensus on a definition of social navigation or how to achieve it. Indeed, at the Social Navigation Symposium [^1] https://sites.google.com/view/socialnavigationsymposium/home, a diverse spectrum of researchers presented a variety of views on robotic social navigation and their approaches to it, including a range of definitions, variants, problems and subproblems.
Ideas presented at the the Symposium included various methods and procedures to evaluate social navigation performance, often involving different experimental setups, evaluation metrics, robot simulators, social datasets, and deployment environments. As the researchers continued their discussions during and after the symposium, a taxonomy of aspects of social navigation began to emerge, which helped clarify the social robot navigation problem and converged to a set of general recommendations on how to evaluate solutions in ways that were more comparable.
This paper summarizes the work done by symposium participants to define the social robot navigation problem, identify a taxonomy of its important aspects, create guidelines for its evaluation, and improve existing simulators by providing a common implementation of these guidelines. After a review of related work in Section II, Section III proposes a definition of social navigation and a strategy for achieving it by respecting other agents’ goals through principles of safety, comfort, legibility, politeness, social competency, agent understanding, proactivity, and responsiveness to context. Section IV reviews the different scientific questions asked by social navigation researchers, and Section V outlines our taxonomy for analyzing social navigation benchmarks, datasets and simulators. Section VI then discusses the metrics that have been developed for measuring social navigation, including subjective human evaluation metrics, analytic metrics that can be directly computed, and research efforts designed to create learned metrics. Section VII discusses the typical scenarios used in social navigation, and Section VIII describes benchmarks built on these scenarios, while Section IX reviews datasets collected on social navigation. Section X reviews simulators and presents our work to create a unified interface across simulators.
@@ II Related Work
The field of social robot navigation is vast and we will not attempt to summarize it here; instead, we refer to multiple recent surveys on social navigation, including [1, 2, 3, 4, 5, 6, 7, 8, 9, 10]. Among recent surveys, [6] specifically focuses on evaluating social robot navigation algorithms, reviewing 177 recent papers to gather evaluation methods, scenarios, datasets, and metrics, using their findings to discuss shortcomings of existing research and to make recommendations for future research directions. Another recent survey by Mavrogiannis et al. [7] focuses on the core challenges of social navigation with respect to navigation algorithms, human behavior models, and evaluation. Our work builds on the works of [6] and [7] and similar surveys to map the field. We contribute a crisp definition of social robot navigation based on discussions held at the 2022 Social Navigation Symposium, an overview of methodologies for research, and a taxonomy of the field which we use to examine existing metrics, scenarios, benchmarks, datasets, and simulators, and shared principles to make social navigation evaluations comparable across the community.
[10] proposes new metrics specifically evaluating the principles defined in [1], comfort, naturalness, and sociability. Our effort proposes to expand the principles in [1] to a broader set including safety, comfort, legibility, politeness, social norms, agent understanding, proactivity, and contextual appropriateness, and proposes a lifecycle of social navigation with recommendations for metrics, scenarios, benchmarks, datasets and simulators, along with specific guidelines for metric usage and recommended metrics.
Beyond social navigation, clear, repeatable, and accessible benchmarks have accelerated progress in fields like computer vision [11] and natural language processing [12, 13, 14] enabling researchers to compare algorithms, revealing limitations of existing solutions, and illuminating promising new directions. Our effort builds on benchmark challenges in traditional robot navigation [15, 16, 17, 18, 19, 20], social navigation benchmarks and challenges [21, 22, 23, 24, 25, 26, 27, 28, 29] and social navigation scenario development [30, 31, 32]. We review social navigation scenarios and benchmarks in depth in Sections VII and VIII; this work contributes guidelines for scenario development, a review of scenarios in the literature, a social navigation scenario card, as well as guidelines for social navigation benchmarking and social navigation dataset development.
@@ III Towards a Definition of Social Navigation
Social navigation has referred to a range of behaviors from simple navigation around dynamic obstacles, to complying with complex social norms, up to navigating with communicative intent. As such, social navigation risks becoming a “suitcase word”, defined by Marvin Minsky [33] as words that carry other concepts inside them, like memory, emotions, or consciousness; these terms must be unpacked to fully understand their meanings.
In this section, we unpack the term “social robot navigation”. First, we briefly examine social robotics and what it means; then we focus more narrowly on the social navigation problem itself. Then we expand this further by examining the subproblems important in social navigation and how context can affect what navigation behaviors are considered social, formulating these subproblems as principles to guide research.
@@ III-A What is a Social Robot?
In their review of Human-Robot Interaction (HRI) for social robotics, Kanda and Ishiguro [34] argue that in addition to navigation (moving robots from place to place) and manipulation (changing objects in the environment) capable robots must also leverage social interactions, i.e., be able to interplay with humans or other robots in the environment to perform tasks. Further, they distinguish robots that simply encounter humans from those which have specifically designed socially interactive features, such as voices, expressive faces or the ability to gesture.
But simply having socially interactive features in a robot does not mean that the quality of its interactions would be acceptable to humans or efficient for other robots; additional principles are needed to apply these features in a positive way. Developing solutions that create high-quality social interactions autonomously is difficult; many social interactions that Kanda and Ishiguro studied could not be accomplished with the technology of the time and required a human to teleoperate the robot. What distinguishes “social” robotics from pure interactivity?
To define “social” more precisely, we examined the terms social and antisocial for humans. Social sometimes means participating in society, i.e., participating in an interacting group whose individuals modify their behavior to accommodate the needs of others while achieving their own. But social has a second meaning: a social individual has outstanding skills to work with others, based on an understanding of their feelings and needs and adapting to them. Antisocial individuals, in contrast, fail to follow the customs of society or live without consideration for others.
Inspired by these terms when applied to humans, we generalize this notion to other agents, and offer this definition of social robot navigation:
@@
This social quality may be reflected through overt behavior changes, such as respecting social norms, or through an understanding of other agents’ needs, feelings, and/or communicating capabilities.
@@ IV Research Methodologies of Social Navigation
Benchmarks require measures and an evaluation methodology for comparing social robot navigation systems. Because social robot navigation is concerned with methods for controlling mobile robots to operate effectively around people, many at the symposium argued social robot evaluation methodologies should involve the collection of human perceptions of robot behavior; however, others argued there are important scientific questions that can be answered, such as an algorithm’s response to dynamic obstacles, which can affect social principles like safety and comfort. Ultimately, the scientific questions a benchmark asks are important in deciding its scope and content. These scientific questions lead to varied types of studies which gather different types of data, which in turn guide the development of methods, creating a lifecycle of social navigation research.
@@ IV-A Research Questions of Social Navigation
The overarching research question of social robot navigation is developing a scientific understanding of the problem sufficient to build computational models that enable robots to perform acceptably in human environments. This involves understanding the factors that influence social navigation, developing models of those factors, and implementing algorithms that take those factors into account. To fairly evaluate how those algorithms compare to each other, we need benchmarks that help us understand their differences and, ideally, to identify the ones that are better. Given the complexity of social navigation, different benchmarks often focus on different aspects of the problem and thus different, more specific scientific questions. Some of these questions arise from traditional robot navigation research and can arguably be evaluated using traditional methods, with adjustments for human participants:
  • How do methods compare with each other against baselines? Some aspects of method evaluation involve quantitative metrics that can be performed in simulation, such as revealing problems in a robot’s safety layer as it faces increasing obstacle densities. However, when human evaluations are required, these are typically conducted in the real world, though toolkits are now coming into use that enable labeling simulated trajectories as well [58].
@@
Success at task performance is often measurable quantitative, but determining whether a robot is satisfying the principles of social navigation is trickier. While the physical aspects of principle P1, safety, may arguably be measureable quantitiatively (at least in the sense that the lack of safety can be measured through damage and collisions), other principles like P4, politeness, often require human evaluation, and still others like P2, comfort, are often explicitly defined in terms of human reactions. Scientific questions involving these subjective aspects therefore generally require measuring human perceptions and reactions to robot behaviors, and are best investigated through HRI studies:
@@
Many benchmarks focus on a subset of the questions above, which is legitimate as different researchers have different aims and different groups have different needs. The consequence, however, is that social navigation evaluation methodologies have become fragmented and a comprehensive evaluation methodology does not yet exist.
Because different lines of research have different needs, we do not aim to provide a specific, singular evaluation protocol to be used for all social navigation methods, but a methodology by which researchers can make principled decisions to guide their own evaluation protocol. Such a protocol will allow researchers to compare social navigation methods along the dimension relevant to a specific subdomain. The direct comparison of methods, via principled evaluation protocols, will allow us to interrogate the strengths and weaknesses of the methods in question and push knowledge discovery in the field and increase method performance.
@@ IV-B Types of Social Navigation Studies
    Robot Deployments: Robot deployments are conducted in environments partially under the control of the experimenters, such as an office, a classroom building, or factory. In this case, robot deployments necessarily involve experimenters informing participants about the robots, which may change their responses compared to someone encountering a robot in the wild; furthermore, participants necessarily develop experience about the robots that can distort human-robot interactions. Symposium participants reported that users unfamiliar to robots were less accepting of errors than robot researchers, who in turn were less accepting than experienced “robot wranglers” responsible for managing the deployment; these anecdotal reports mirror studies which found evidence that both general computer user skill level [61] and familiarity with particular robots [62, 63, 64] could improve assessments of robot capabilities and behavior. Semi-controlled robot deployments are similar to, but less naturalistic than field studies, but because robot deployment environments are more controlled than true in-the-wild studies, larger scale is often more practicable by conducting experiments over a longer period of time. For example, [65] collected 1,000 kilometers of navigation data in a set of buildings on a college campus, and the system described in [31] was part of a deployment in two Google buildings that collected over 3,000 kilometers of data.
@@ IV-C Lifecycle of Social Navigation Research
    Data Collection: Field studies and robot deployments can be used for the first step of the scientific process: data collection. Ideally, field studies and robot deployments should be used for more than just A/B testing; they should be used to generate datasets that can be shared to extend the power of the social navigation research community to collect data at scale. Data that can be collected includes but is not limited to robot and human behavior, surveys (e.g., subject’s opinions on safety or comfort), or biometric data (e.g., heart rate, skin impedance).
@@
    Laboratory Experiments: Many scientific questions about human-robot interaction can be conducted even if large-scale datasets or issue statistics do not exist. Research groups not positioned to conduct large-scale studies or deployments can nevertheless formulate scientific questions and answer them. Where feasible, these experiments should use benchmarking procedures and metrics which have been validated by the research community, such as those discussed in Section VI. Ideally, these should also use scenarios identified as frequent issues in the target domain.
@@ IV-D Guidelines for Real-world Studies
Guideline R3: Ensure Clear Scientific Objectives. As real-world experiments are valuable, the purpose of real-world benchmark studies should be clearly defined and the use of the data specified. Conducting an experiment which costs time and money and exposes humans, robots and their environments to risk should be justified with a clear notion of what is to be learned from conducting the experiment.
@@ V A Taxonomy of Social Navigation
Our workgroup does not aim to provide a single comprehensive evaluation protocol to be used for all social navigation research, because social navigation research spans a wide range of domains with different goals and priorities. Instead, we are developing principles and guidelines to guide the development of protocols that enable fair comparisons of methods. To enable this comparison, we have developed a taxonomy for social navigation research in terms of the metrics, datasets, simulators, and benchmarks used, analyzed with a common vocabulary for factors of analysis.
@@ V-A A Taxonomy for Analysis
  • Metrics: Recent surveys of social navigation metrics have uncovered close to a hundred different metrics in use (see  [6, 8] for recent reviews). However, these metrics can be grouped by broader features, such as algorithmically computed (such as Success-weighted Path Length (SPL) [15]) or gathered by surveys of humans (such as comfort ratings). Surveyed metrics can further be divided into those collected explicitly via questionnaires (e.g., about perceived comfort) or implicitly through sensors (such as stress in human facial expressions). Algorithmic metrics in turn can be hand-crafted or learned from data gathered from surveys. Other axes of metrics include the type of variable(s) being modeled, and whether metrics cover the behavior of the robot at a specific point in time (step wise) or during a whole navigation task (episode wise). Metrics are further discussed in Section VI.
@@ VI-A1 Taxonomy based on their nature
Surveyed metrics are usually human ratings of desired properties of social robot navigation, e.g. safety, comfort, legibility. They can in turn be classified into questionnaire-based (in-situ or ex-situ), where the rates are explicitly requested, or sensor-based, where the rates are transduced from sensor data. Although they arguably are the best (or correct) metrics to measure success for social navigation, they are expensive, difficult to scale, and time-consuming. While small-scale human studies are commonly conducted, the results can have high variance and can be non-reproducible. To address this shortcoming, researchers have also created a variety of algorithmic metrics that serve as proxies for surveyed metrics. These algorithmic metrics are cheap to compute and reproducible, two properties that are key for benchmarking. Unlike in traditional navigation, where SPL [15] is a commonly accepted metric, social navigation has no single metric of reference. Instead, method comparison is usually performed using multiple metrics.

### v4

@@ Abstract
Figure 1: We identify eight broad principles of social robot navigation - including safety, comfort, legibility, politeness, social competency, agent understanding, proactivity, and contextual appropriateness - which motivate specific guidelines for experiments, metrics, scenarios, benchmarks, datasets, and simulators. Principles and guidelines are labeled with two-letter codes, with P for principles, R for real-world issues, M for metrics, N for scenarios, B for benchmarks, D for datasets, and S for simulators.
@@ I Introduction
The study of social robot navigation has a long history, but a crisp definition of what makes navigation “social” remains elusive. Researchers on social robot navigation often have a personal sense of what it is and use that intuition to guide their research into how to make robots move better around people, but the field does not yet have a consensus on a definition of social navigation or how to achieve it. Indeed, at the Social Navigation Symposium [^1] https://sites.google.com/view/socialnavigationsymposium/home, a diverse spectrum of researchers presented a variety of views on what robotic social navigation is and their approaches to solving it, including a range of definitions, variants, problems and subproblems.
Ideas presented at the the Symposium included a variety of methods to evaluate social navigation performance, involving different experimental setups, evaluation metrics, robot simulators, social datasets, and deployment environments. As the researchers continued their discussions following the symposium, a taxonomy of aspects of social navigation began to emerge, which helped clarify the social robot navigation problem and converged to a set of general recommendations on how to evaluate solutions in ways that were more comparable.
This paper summarizes our work to define the social robot navigation problem, identify a taxonomy of its important aspects, create guidelines for its evaluation, and define a common API to make evaluations more comparable. After a review of related work in Section II, Section III proposes a definition of social navigation and a strategy for achieving it by respecting other agents’ goals through principles of safety, comfort, legibility, politeness, social competency, agent understanding, proactivity, and responsiveness to context. Section IV reviews the different scientific questions asked by social navigation researchers, and Section V outlines our taxonomy for analyzing social navigation benchmarks, datasets and simulators. Section VI then discusses the metrics that have been developed for measuring social navigation, including subjective human evaluation metrics, analytic metrics that can be directly computed, and research efforts designed to create learned metrics. Section VII discusses the typical scenarios used in social navigation, and Section VIII describes benchmarks built on these scenarios, while Section IX reviews datasets collected on social navigation. Section X reviews simulators and presents our work to create a unified interface across simulators.
@@ II Related Work
The field of social robot navigation is vast and we will not attempt to summarize it; instead, we refer to many recent surveys on social navigation [1, 2, 3, 4, 5, 6, 7, 8, 9, 10]. Among these, [6] focuses on evaluating social robot navigation algorithms, reviewing 177 recent papers to gather evaluation methods, scenarios, datasets, and metrics, using their findings to discuss shortcomings of existing research and to make recommendations for future research directions. Another recent survey by Mavrogiannis et al. [7] focuses on the core challenges of social navigation with respect to navigation algorithms, human behavior models, and evaluation. Our work builds on the works of [6] and [7] and similar surveys to map the field. We contribute a crisp definition of social robot navigation based on discussions held at the 2022 Social Navigation Symposium, an overview of methodologies for research, and a taxonomy of the field which we use to examine existing metrics, scenarios, benchmarks, datasets, and simulators, and shared principles to make social navigation evaluations comparable across the community.
[10] proposes new metrics evaluating the principles defined in [1], comfort, naturalness, and sociability. We expand the principles in [1] to a broader set including safety, comfort, legibility, politeness, social norms, agent understanding, proactivity, and contextual appropriateness, and proposes a lifecycle of social navigation with recommendations for metrics, scenarios, benchmarks, datasets and simulators, along with guidelines for metric usage and recommended metrics.
Beyond social navigation, clear, repeatable, and accessible benchmarks have accelerated progress in fields like computer vision [11] and natural language processing [12, 13, 14] enabling researchers to compare algorithms, revealing limitations of existing solutions, and illuminating promising new directions. Our effort builds on benchmark challenges in traditional robot navigation [15, 16, 17, 18, 19, 20], social navigation benchmarks and challenges [21, 22, 23, 24, 25, 26, 27, 28, 29] and social navigation scenario development [30, 31, 32]. We review social navigation scenarios and benchmarks in Sections VII and VIII. We contribute guidelines for scenario development, a review of scenarios in the literature, a social navigation scenario card, as well as guidelines for social navigation benchmarking and dataset development.
@@ III Towards a Definition of Social Navigation
Social navigation refers to a range of behaviors from simple navigation around dynamic obstacles, to complying with complex social norms, up to navigating with communicative intent. As such, it risks becoming a “suitcase word”, defined by Marvin Minsky [33] as words that carry other concepts inside them, like memory, emotions, or consciousness; these terms must be unpacked to fully understand their meanings.
In this section, we unpack the term “social robot navigation”. First, we examine social robotics and what it means; then we focus on the social navigation problem itself. Then we examine the subproblems of social navigation and how context can affect what behaviors are considered social, formulating these subproblems as principles to guide research.
@@ III-A What is a Social Robot?
In their review of Human-Robot Interaction (HRI) for social robotics, Kanda and Ishiguro [34] argue that in addition to navigation (moving robots from place to place) and manipulation (changing objects in the environment) capable robots must also leverage social interactions, i.e., be able to interplay with humans or other robots to perform tasks. Further, they distinguish robots that simply encounter humans from those which have specifically designed socially interactive features, such as voices, expressive faces or the ability to gesture.
But simply having socially interactive features in a robot does not mean that the quality of its interactions would be acceptable to humans or efficient for other robots; additional principles are needed to apply these features in a positive way. Developing solutions that create high-quality social interactions autonomously is difficult; many social interactions that Kanda and Ishiguro studied were beyond the technology of the time and required a human to teleoperate the robot. What distinguishes “social” robotics from pure interactivity?
To define “social” more precisely, we examined the terms social and antisocial for humans. Social sometimes means participating in society, i.e., participating in an interacting group whose individuals modify their behavior to accommodate the needs of others while achieving their own. But social has a second meaning: a social individual has outstanding skills to work with others, based on an understanding of their feelings and needs and adapting to them. Antisocial individuals, in contrast, fail to follow the customs of society or live without consideration for others. Inspired by these terms when applied to humans, we generalize this notion to other agents, and offer this definition of social robot navigation:
@@
This social quality may be reflected through overt behavior changes, such as respecting social norms, or through understanding other agents’ needs, feelings, and capabilities.
@@ IV Research Methodologies of Social Navigation
Benchmarks require measures and an evaluation methodology for comparing social robot navigation systems. Because social navigation is concerned with methods for controlling mobile robots to operate effectively around people, many at the symposium argued social robot evaluation methodologies should involve the collection of human perceptions of robot behavior; however, others argued there are important scientific questions, such as an algorithm’s response to dynamic obstacles, which can affect social principles like safety and comfort. Ultimately, the scientific questions a benchmark asks are important in deciding its scope and content. These scientific questions lead to varied types of studies which gather different types of data, which in turn guide the development of methods, creating a lifecycle of social navigation research.
@@ IV-A Research Questions of Social Navigation
The overarching research question of social robot navigation is developing a scientific understanding of the problem sufficient to build computational models that enable robots to perform acceptably in human environments. This involves understanding the factors that influence social navigation, developing models of those factors, and implementing algorithms that take those factors into account. To fairly evaluate how algorithms compare to each other, we need benchmarks that help us understand their differences and identify which ones are better. Given the complexity of social navigation, different benchmarks often focus on different aspects of the problem and thus different, more specific scientific questions. Some of these questions arise from traditional robot navigation research and can arguably be evaluated using traditional methods, with adjustments for human participants:
  • How do methods compare with each other against baselines? Some aspects of method evaluation involve quantitative metrics measurable in simulation, such as revealing problems in a robot’s safety layer as it faces increasing obstacle densities. However, when human evaluations are required, these are typically conducted in the real world, though toolkits are now coming into use that enable labeling simulated trajectories as well [58].
@@
Success at task performance is often measurable quantitative, but determining whether a robot is satisfying the principles of social navigation is trickier. While the physical aspects of Principle P1, Safety, may arguably be measureable quantitiatively (at least in the sense that the lack of safety can be measured through damage and collisions), others like Principle P4, Politeness, often require human evaluation, and still others like Principle P2, Comfort, are often explicitly defined in terms of human reactions. Scientific questions involving these subjective aspects therefore generally require measuring human perceptions and reactions to robot behaviors, and are best investigated through HRI studies:
@@
Many benchmarks focus on a subset of these questions because different researchers have different aims and different groups have different needs. As a result, social navigation evaluation methodologies have become fragmented and a comprehensive evaluation methodology does not yet exist.
Because different lines of research have different needs, we do not aim to provide one evaluation protocol for all social navigation methods, but a methodology by which researchers can make principled decisions to guide their own evaluations. Such a protocol will allow researchers to compare social navigation methods along the dimension relevant to a specific subdomain. The direct comparison of methods, via principled evaluation protocols, will allow us to interrogate the strengths and weaknesses of methods, push knowledge discovery in the field, and increase method performance.
@@ IV-B Types of Social Navigation Studies
    Robot Deployments: Robot deployments are conducted in environments partially under the control of the experimenters, such as an office, a classroom building, or factory. In this case, robot deployments necessarily involve experimenters informing participants about the robots, which may change their responses compared to someone encountering a robot in the wild; furthermore, participants necessarily develop experience about the robots that can distort human-robot interactions. Symposium participants reported that users unfamiliar to robots were less accepting of errors than robot researchers, who in turn were less accepting than experienced “robot wranglers” responsible for managing the deployment; these anecdotal reports mirror studies which found evidence that both general computer user skill level [61] and familiarity with particular robots [62, 63, 64] could improve assessments of robot capabilities and behavior. Semi-controlled robot deployments are similar to, but less naturalistic than field studies, but because robot deployment environments are more controlled than true in-the-wild studies, larger scale is often more practicable by conducting experiments over a longer period of time. For example, [65] collected 1,000 kilometers of indoor navigation on a college campus, and the system described in [31] was part of a deployment at Google that collected over 3,000 kilometers of data.
@@ IV-C Lifecycle of Social Navigation Research
    Data Collection: Field studies, robot deployments and staged interactions can be used for the first step of the scientific process: data collection. Ideally, these should be used for more than just A/B testing; they should be used to generate datasets that can be shared to extend the power of the social navigation research community to collect data at scale. Data that can be collected includes but is not limited to robot and human behavior, surveys (e.g., subject’s opinions on safety or comfort), or biometric data (e.g., heart rate, skin impedance).
@@
    Laboratory Experiments: Many scientific questions about human-robot interaction can be conducted even if large-scale datasets or issue statistics do not exist. Research groups not able to conduct large-scale studies or deployments can nevertheless formulate scientific questions and answer them. Where feasible, these experiments should use benchmarking procedures and metrics validated by the research community, such as those discussed in Section VI. Ideally, these should use scenarios identified as frequent issues in the target domain.
@@ IV-D Guidelines for Real-world Studies
Guideline R3: Ensure Clear Scientific Objectives. As real-world experiments are expensive, the purpose of real-world benchmark studies should be clearly defined and the use of the data specified. Conducting an experiment which costs time and money and exposes humans, robots and their environments to risk should be justified with a clear notion of what is to be learned from conducting the experiment.
@@ V A Taxonomy of Social Navigation
Creating principles and guidelines that are broadly useful to the social navigation community requires understanding the research efforts already underway. Therefore, we have developed a taxonomy for social navigation research in terms of the metrics, datasets, simulators, and benchmarks used, analyzed with a common vocabulary for factors of analysis.
@@ V-A A Taxonomy for Analysis
  • Metrics: Social navigation metrics enable us to quantify the performance of algorithms. Recent surveys of social navigation metrics have uncovered close to a hundred different metrics in use (see  [6, 8] for recent reviews). Some metrics are algorithmically computed, while others are gathered by surveying humans, either explicitly via questionnaires or implicitly through sensors measuring affect. Algorithmic metrics in turn can be hand-crafted or learned from data gathered from surveys. Other axes of metrics include the type of variable(s) being modeled, and whether metrics cover the behavior of the robot at a specific point in time (step wise) or during a whole navigation task (episode wise). The nuances of metrics are discussed in depth Section VI.
@@ VI-A1 Taxonomy based on their nature
Surveyed metrics are usually human ratings of desired properties of social robot navigation, e.g. safety, comfort, or legibility. They can be classified into questionnaire-based (in-situ or ex-situ), where the ratings are explicitly requested, or sensor-based, where the ratings are transduced from sensor data. Although surveyed metrics are (arguably) the best way to measure social navigation success, they are expensive, difficult to scale, and time-consuming. While small-scale human studies are commonly conducted, results can have high variance and can be non-reproducible. To address this shortcoming, researchers have also created a variety of algorithmic metrics that serve as proxies for surveyed metrics. These algorithmic metrics are cheap to compute and reproducible, properties that are key for benchmarking. Unlike traditional navigation, where SPL [15] is a commonly accepted metric, social navigation has no single metric of reference. Instead, method comparison is usually performed using multiple metrics.